OpenStudy (anonymous):

How do i find the average speed for his question: "A bicycle racer rides from a starting marker at 10m/s. She then rides back along the same rout from the turnaround marker at 16m/s. What is her average speed for the whole race?"

OpenStudy (anonymous):

well my book says you have to use speen and velocity

OpenStudy (anonymous):

I get that, (y2-y1/x2-x1) but theres no info in the question about the actual amount of distance traveled or actual time it took to travel

OpenStudy (anonymous):

You may be able to add both velocities up and divide them up by how many there are so, \[\frac{ v_1+v_2 }{ 2 }\]

OpenStudy (anonymous):

I thought of that, but the book im using said that that is incorrect and the answer is actually 12m/s, not 13m/s like you and i thought

OpenStudy (anonymous):

Find all the distances. Average Speed is ALWAYS total distance/total time.

OpenStudy (anonymous):

Assuming you have the times taken. Is that the whole question?

OpenStudy (anonymous):

Yeah, that's the entirety of the question.

OpenStudy (anonymous):

We can substitute values around to come up with in that equation to give, \[\frac{ 2v_1+v_2 }{ (v_1+v_2) }\]

OpenStudy (anonymous):

but that gives 1.384, not 12m/s as the book's answer states

OpenStudy (anonymous):

Sorry the addition sign is not suppose to be there its all multiplied at the top

OpenStudy (anonymous):

\[\frac{ 2v_1v_2 }{ (v_1+v_2) }\]

OpenStudy (anonymous):

Eureka! Thank you!

OpenStudy (anonymous):

My pleasure

OpenStudy (anonymous):

Good work, John.

OpenStudy (anonymous):

I believe that solves your question! Please mark the question as Closed. Thanks :)